OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x + 3y = -1 6x - 5y = 16

OpenStudy (anonymous):

what subject is this?

jhonyy9 (jhonyy9):

5x+3y= -1 * 5 6x-5y= 16 * 6 ----------------- 30x +15y = -5 30x -30y = 96 ---------------- subtract second from first 0 45y = -101 y = - 101 / 45